# Causal Climate Emulation with Bayesian Filtering

### v1

@@ Appendix G Annual temperatures ranges generated by PICABU
Here we illustrate the average intra-annual temperature range for NorESM2 and PICABU simulations over a 100-year simulation. For each grid cell, for each year, we take the month with the maximum value, and the month with the minimum value, and then compute the difference to give the intra-annual range. This quantity is then averaged over the 100 years of simulation for both NorESM2 and PICABU to generate Figure 13. The spatial distribution of the ranges is reasonable, with some notable exceptions over land (such as sub-Saharan Africa), where the range is much too small. The ranges over the ocean largely agree.
@@ Appendix H Indian Ocean Dipole and Atlantic Multidecadal Oscillation emulation
Figure 14 shows example time series of NorESM’s simulated IOD and AMO, alongside PICABU’s emulation, and also shows the power spectra of the time series of both modes of variability. The quantitative performance is given in Table 4.
@@ Appendix I Results of CESM2-FV2 emulation
In addition to training and evaluating PICABU for NorESM2 emulation, we also separately trained and evaluated on a second climate model, CESM2-FV2. We did this to investigate whether our model is able to emulate another climate model, how sensitive PICABU’s training is to hyperparameters, and whether our findings regarding PICABU’s performance are consistent for another climate model. It should be noted that NorESM2 is based on CESM2, and therefore has many shared model components. NorESM2 has a different model for the ocean, ocean biogeochemistry, and atmospheric aerosols. It also includes specific modifications and tunings of the dynamics of the atmosphere model [Seland et al., 2020]. However, the two models are not completely independent.
@@
In Figure 15 we illustrate example time series of global mean surface temperature (GMST), Niño3.4, the IOD, and the AMO, as also shown for NorESM2, giving a qualitative illustration of the variability generated by PICABU for each. Figure 16 and Figure 17 show the power spectra of PICABU and the ground-truth CESM2 model for the different indices, showing the spectra on a standard and log plot, respectively. The quantitative results are presented in Table 5.
@@
Figure 15: For CESM2-FV2, similar to NorESM2, PICABU generates realistic ENSO and IOD variability, but does not capture GMST short-term variability. We provide example time series of the Niño3.4 index and GMST for both the ground-truth climate model CESM2-FV2 and PICABU over 50 years. Example time series of the IOD and AMO are also shown.
@@
Figure 17: The power spectra on a log scale for four measures of climate variability, ENSO, IOD, GMST and AMO are shown, both for the ground truth CESM2-FV2 model, and rollouts generated by PICABU.
@@
Appendix J Ablation of number of latents and sparsity coefficient
We investigate the performance of PICABU when we vary the number of latent variables. Figure 18 shows the clustering for different numbers of latent variables. Across different numbers of latents, PICABU learns similar clustering, with larger clusters being progressively split up as the number of latents increases, while some prominent clusters remain for most numbers of latents. Figure 19 shows the power spectra for Niño3.4 and GMST for models trained with different numbers of latents. All models are stable, though notably for small numbers of latents, PICABU does not represent the 3-year peak in the ENSO power spectrum.
Figure 18: Spatial aggregation learned by PICABU for different numbers of latent variables. The spatial aggregation learned by PICABU shows variation as the number of latents is varied. However, some clusters are reasonably consistent across the different number of latents.
@@
Appendix K Multivariable emulation
While most results in this work focus on a single variable, skin temperature, we illustrate that multivariable emulation is possible. PICABU can be extended to include multiple variables by learning latents for each individual variable, and allowing all of these latents to influence each other, including across variables (this is increasingly expensive if the number of variables and the number of latents is large). Figure 21 shows two previous timesteps of data, the reconstruction from the model (simply encoding and decoding the given timestep), and the next timestep prediction.
Figure 21: PICABU may be used for multivariable emulation. Each row shows a different variable, starting from the top: 1) Skin temperature, 2) surface temperature, 3) sea-level pressure, and 4) precipitation. The first column shows the ground truth at $t-1$ (previous timestep), and the second column shows the ground truth at time $t$. The third column shows the reconstruction of each variable, and the fourth column shows PICABU’s prediction at time $t$.
Appendix L Training objective for non-causal model
In Section 4.3, the following objectives are optimized when training the model without sparsity (Equation 14) and without orthogonality (Equation 15), respectively:
@@
Appendix M Experiments compute resources
To train PICABU on 800 years of temperature data, we have used the following compute resources: 2 RTX8000 GPUs, with 48GB of RAM each, for $\approx$10 hours. These resources scale linearly with the number of variables, number of latents, or number of input timesteps $\tau$. The synthetic experiments are much faster ($\approx$0.5 hour on 1 RTX8000 GPU), as it is lower dimensional. All experiments were run on an internal cluster. We performed experiments related to hyperparameter tuning on these resources (Appendix N).
Appendix N Hyperparameter values
We provide the values of the hyperparameters that we use for the model training and for the Bayesian filtering in Table 6. These hyperparameters were determined with considerable manual tuning, with the values of the spectral penalties being important for model performance.
@@
Appendix O Licenses for existing assets
The datasets used in this work are CMIP6, associated with the Creative Commons Attribution 4.0 International license (CC BY 4.0), CESM, provided by UCAR. This work builds upon ClimateSet, published in NeurIPS 2023 and publicly available on arXiv, CDSD, publicly available on arXiv, and SAVAR, associated with the CC BY 4.0 license.

### v2

@@ Appendix G Annual temperatures ranges generated by PICABU
Here we illustrate the average intra-annual temperature range for NorESM2 and PICABU simulations over a 100-year simulation. For each grid cell, for each year, we take the month with the maximum value, and the month with the minimum value, and then compute the difference to give the intra-annual range. This quantity is then averaged over the 100 years of simulation for both NorESM2 and PICABU to generate Figure˜14. The spatial distribution of the ranges is reasonable, with some notable exceptions over land (such as sub-Saharan Africa), where the range is much too small. The ranges over the ocean largely agree.
@@ Appendix H Indian Ocean Dipole and Atlantic Multidecadal Oscillation emulation
Figure˜15 shows example time series of NorESM’s simulated IOD and AMO, alongside PICABU’s emulation, and also shows the power spectra of the time series of both modes of variability. The quantitative performance is given in Table˜4.
@@ Appendix I Results of CESM2-FV2 emulation
In addition to training and evaluating PICABU for NorESM2 emulation, we also separately trained and evaluated on a second climate model, CESM2-FV2. We did this to investigate whether our model is able to emulate another climate model, how sensitive PICABU’s training is to hyperparameters, and whether our findings regarding PICABU’s performance are consistent for another climate model. It should be noted that NorESM2 is based on CESM2 and, therefore, has many shared model components. NorESM2 has a different model for the ocean, ocean biogeochemistry, and atmospheric aerosols. It also includes specific modifications and tunings of the dynamics of the atmosphere model [Seland et al., 2020]. However, the two models are not completely independent.
@@
In Figure˜16 we illustrate example time series of global mean surface temperature (GMST), Niño3.4, the IOD, and the AMO, as also shown for NorESM2, giving a qualitative illustration of the variability generated by PICABU for each. Figure˜17 and Figure˜18 show the power spectra of PICABU and the ground truth CESM2 model for the different indices, showing the spectra on a standard and log plot, respectively. The quantitative results are presented in Table˜5.
@@
Figure 16: For CESM2-FV2, similar to NorESM2, PICABU generates realistic ENSO and IOD variability, but does not capture GMST short-term variability. We provide example time series of the Niño3.4 index and GMST for both the ground truth climate model CESM2-FV2 and PICABU over 50 years. Example time series of the IOD and AMO are also shown.
@@
Figure 18: PICABU predictions exhibit too much power at high frequencies. The power spectra on a log scale for four measures of climate variability, ENSO, IOD, GMST, and AMO are shown, both for the ground truth CESM2-FV2 model, and rollouts generated by PICABU.
@@
Appendix J Choice of constant variance in the Bayesian filter
When running a standard Bayesian filter, the variance for the distribution of the predicted samples is typically the variance of the distribution observed in the model. Table˜2 shows that PICABU learns a distribution with higher variance than the observations. To avoid propagating this higher uncertainty through the prediction, we instead use the assumption of constant variance, which is well-suited to the pre-industrial control data, and estimate this constant variance from the observations directly. Table˜6 shows summary statistics when using constant variance vs. the model’s inferred variance.
Mean Std. Dev. Range LSD Mean Std. Dev. Range LSD
GMST Niño3.4
Ground truth 0 0.177 1.357 0 0 0.927 5.73 0
Constant var. -0.0518 0.277 2.02 0.444 -0.0914 1.05 5.99 0.191
Inferred var. -0.0742 0.324 2.15 0.464 -0.198 1.27 7.70 0.206
IOD AMO
Ground truth 0 0.877 7.36 0 0 0.368 2.58 0
Constant var. 0.186 0.753 5.36 0.0772 0.0538 0.445 3.27 0.248
Inferred var. 0.110 0.845 5.46 0.0753 0.0157 0.498 3.27 0.261
Table 6: Using the model’s estimated variance in the Bayesian filter slightly degrades the predictions. The table shows statistics (mean, standard deviation, range, log-spectral distance) for the ground truth, prediction using a constant variance estimated from the observations, and the model’s inferred variance. The predictions degrade slightly when using the inferred variance, especially the standard deviation and range, as PICABU tends to overestimate the variance of the data.
Appendix K Ablation of number of latents and sparsity coefficient
We investigate the performance of PICABU when we vary the number of latent variables. Figure˜19 shows the clustering for different numbers of latent variables. Across different numbers of latents, PICABU learns similar clustering, with larger clusters being progressively split up as the number of latents increases, while some prominent clusters remain for most numbers of latents. Figure˜20 shows the power spectra for Niño3.4 and GMST for models trained with different numbers of latents. All models are stable, though notably for small numbers of latents, PICABU does not represent the 3-year peak in the ENSO power spectrum.
Figure 19: Spatial aggregation learned by PICABU for different numbers of latent variables. The spatial aggregation learned by PICABU shows variation as the number of latents is varied. However, some clusters are reasonably consistent across the different numbers of latents.
@@
Appendix L Multivariable emulation
While most results in this work focus on a single variable, skin temperature, we illustrate that multivariable emulation is possible. PICABU can be extended to include multiple variables by learning latents for each individual variable, and allowing all of these latents to influence each other, including across variables (this is increasingly expensive if the number of variables and the number of latents is large). Figure˜22 shows two previous timesteps of data, the reconstruction from the model (simply encoding and decoding the given timestep), and the next timestep prediction.
Figure 22: PICABU may be used for multivariable emulation. Each row shows a different variable, starting from the top: 1) Skin temperature, 2) surface temperature, 3) sea-level pressure, and 4) precipitation. The first column shows the ground truth at t1t-1 (previous timestep), and the second column shows the ground truth at time $t$. The third column shows the reconstruction of each variable, and the fourth column shows PICABU’s prediction at time $t$.
Appendix M Training objective for non-causal model
In Section˜4.3, the following objectives are optimized when training the model without sparsity (Equation˜14) and without orthogonality (Equation˜15), respectively:
@@
Appendix N Experiments compute resources
To train PICABU on 800 years of climate model data, we used the following compute resources: 2 RTX8000 GPUs, with 48GB of RAM each, for $\approx$10 hours. These resources scale linearly with the number of variables, number of latents, or number of input timesteps $\tau$. The synthetic experiments are much faster ($\approx$0.5 hour on 1 RTX8000 GPU), as it is lower dimensional. For our autoregressive rollouts emulating climate model data, the use of the Bayesian filter introduces some computational overhead compared to direct deterministic rollouts. However, the operations required for the Bayesian filter (sampling, decoding, and fast Fourier transforms) can be batched at each step and performed rapidly. In general, we are able to simulate 100 years (1200 months) in less than 6 minutes on a single RTX8000 GPU with 48GB memory, with $N=300$ and $R=10$ (with the computational time scaling linearly with $N\cdot R$).
All experiments were run on an internal cluster. We conducted hyperparameter tuning on these resources (Appendix˜O).
Appendix O Hyperparameter values
We provide the values of the hyperparameters that we use for the model training and for the Bayesian filtering in Table˜7. These hyperparameters were determined with manual tuning, with the values of the spectral penalties being important for model performance. We performed a search over the  20 parameters described in Table˜7, with  100 runs of PICABU.
@@
After the initial set of parameters was selected on NorESM2, we performed minimal parameter tuning to train PICABU on CESM2, as described in Appendix˜I, and the hyperparameters performed well on the simulated dataset (SAVAR) where only the sparsity of the final graph was changed. These results suggest that in practice, our hyperparameters will provide a sound baseline for adaptation to other climate models. Moreover, Appendix˜M shows that our default hyperparameters are robust while varying other aspects of the model (e.g., number of latents, sparsity).
@@
Appendix P Observation-space interventions
In addition to intervention in observation space on ENSO, we carried out interventions in observation space to perturb the IOD, to explore whether the model responded to these interventions in a manner consistent with known teleconnections (Figure˜23). We directly increase the temperature value in the region for the Niño 3.4 intervention, while for the IOD we simply increase the temperature of the western Indian Ocean to increase the IOD index. In line with expected teleconnections, increasing El Niño, as measured by the Niño 3.4 index, causes increased global temperatures with a correlation of 0.9. We also observed, in line with Cai et al. [2009], that in the model there is a causal teleconnection between the springtime strength of the Indian Ocean Dipole and temperatures over Australia, with a correlation of 0.87. Furthermore, interventions over Alaska of a similar magnitude had minimal effects on GMST, with a correlation of 0.08 (not shown).
Figure 23: Intervention in the Niño 3.4 region leads to increased GMST, while intervention in the IOD region increases temperatures in Australia. (Left) Intervened temperatures for the Niño 3.4 region (x-axis) are positively correlated with GMST (y-axis). (Right) Intervened temperatures for the IOD region (x-axis) are positively correlated with temperatures over Australia (y-axis).
The maps of these interventions are shown in Figure˜24, where we intervene to increase the temperature anomaly over Niño 3.4 and Alaska by 2 standard deviations, and we increase the value of the IOD index by 0.5.
Figure 24: Maps of the effect of an intervention in observation space where we separately increase the temperature over the Niño 3.4 area and Alaska, and increase the IOD. (Top row) Increased global temperatures are observed for the Niño intervention, showing a large effect in the tropics and a smaller effect outside the tropics. (Middle row) Little difference is observed when intervening on Alaskan temperatures, despite the similarity of the magnitude of the intervention. (Bottom row) The increase in the IOD index leads to an increase in temperatures over Australia and Africa.
Appendix Q Intervention in latent space
We can also directly intervene on latent variables to explore the effect of interventions on latents which correspond to known modes of climate variability. In Figure˜25, which illustrates the effect of intervening on the latent variable that most closely overlaps with the region of the Pacific used to define the Niño 3.4 index.
Figure 25: Intervention on the latent variable that describes the ENSO state. The left panel shows mapping from latents to observations (colored grid points), and the intervened latent variable. We increase its value, which then influences the latents at the next timestep through the learned causal graph. No other latent variables are intervened on. The middle panel shows the original, unintervened next-step prediction, and the right panel shows the intervened prediction.
The single-parent structure allows for a clear correspondence between latents and the physical quantity in the climate model. We verified that the learned Niño 3.4 latent variable is positively correlated with the state of ENSO before performing the counterfactual experiment, by plotting the decoding function from this latent to the corresponding observations (Figure˜26). This yielded a linear relationship with a correlation of 0.98.
Figure 26: Effect of the intervention on latent variable 16 (most overlapping with the Niño3.4 region) on the physical quantity of temperature in the grid cells within the Niño 3.4 region.
Appendix R Licenses for existing assets
The datasets used in this work are CMIP6, associated with the Creative Commons Attribution 4.0 International license (CC BY 4.0), and CESM2, provided by UCAR. This work builds upon ClimateSet, published in NeurIPS 2023 and publicly available on arXiv, CDSD, publicly available on arXiv, and SAVAR, associated with the GPLv3 license.